OpenStudy (anonymous):

Use the Ratio Test to find the radius of convergence of SUM [x^n/2n^2 ]. So far I have...

OpenStudy (amistre64):

take the limit of an/an-1

OpenStudy (amistre64):

\[\lim \frac{x^n}{2n^2}\frac{2(n-1)^2}{x^{n-1}}\]

OpenStudy (amistre64):

simplify and then take out all the notn parts

OpenStudy (amistre64):

\[\lim x\frac{(n-1)^2}{n^2}\] \[|x|\ \lim \frac{(n-1)^2}{n^2}\to\ |x|*1\] |x|<1 you have a radius of 1

OpenStudy (anonymous):

GREAT! I have that!!

OpenStudy (amistre64):

.... if anything, you can always dbl chk with the wolf

OpenStudy (anonymous):

haha, "the wolf"!!

OpenStudy (anonymous):

Question: does it matter if |x|<1 or \[|x|\le 1\] ??

OpenStudy (amistre64):

no, the radius is still 1

OpenStudy (amistre64):

the interval of convergence is -1 to 1

OpenStudy (anonymous):

Oooooh